OpenStudy (anonymous):

Geometry Triangle ABC is similar to Triangle XYZ. What is the length of segment XZ if BC= 9m, AC= 6m and YZ= 3m?

OpenStudy (anonymous):

hahaa. 6m atah yan. =D

OpenStudy (anonymous):

pano mo na-solve?

OpenStudy (anonymous):

haha, diba ung given, BC= 9m, AC= 6m and YZ= 3m , soooo ang magiging line segment na yan is line segment BAC and line segment YXZ. kasi nga mas mahaba ung BC kaysa AC.. tsaka diba similar sila. kaya YXZ, tapos ang size ng AC 6m dibaaa,, dahil similar sila 6m din ung XZ,, gets?? haha

OpenStudy (anonymous):

haha sige thanks :))

OpenStudy (dandandan):

2m

OpenStudy (dandandan):

use the principle of similar triangles. they do not possess the same dimensions.

OpenStudy (anonymous):

if the triangles are similar, then this means ONE thing. that is, there sides are in proportion. furthermore, the CORRESPONDING sides of both triangles differ by some constant number that means XZ/AC = YZ/BC XZ = AC(YZ/BC) make your substitutions XZ = 6(3/9) XZ = 2